# PAMS: The Perseus Arm Molecular Survey – I. Survey description and first results

### v1

@@ 5 Discussion
In Section 4.1, we calculated factors for 12CO and 13CO (3–2) between the Inner and Outer Galaxy by comparing Herschel-derived H2 column densities with data from the COHRS (Dempsey et al., 2013; Park et al., 2023), CHIMPS (Rigby et al., 2016), and PAMS data, summarising the results in Tables 4 and 4. The Inner Galaxy sight-line used was at °, whose emission is dominated by the W43 star-forming region at a distance of 5.2 kpc (Urquhart et al., 2018), corresponding to kpc, while the PAMS regions are located close to kpc.
While we found that the distributions of and values calculated on a pixel-by-pixel basis were significantly different between the two regions (and this is confirmed by both Kolmogorov-Smirnov and Anderson-Darling statistical tests), the differences between the representative values are rather mild. For 12CO (3–2), values reported by all four methods differ by less than 50 per cent. The distribution of pixel values overlap substantially, though the Outer Galaxy distribution has a greater proportion of pixels with higher values, but these originate in relatively cold and low-column density gas and thus contribute relatively little of the CO emission.
The differences are more significant in 13CO (3–2), where the median value of in the Inner Galaxy is a factor of two higher than the equivalent value in the Outer Galaxy, and the distributions of pixel values are more different. However, again the global values (method iv)) agree to within 50 per cent, and the fitted values (method i)) are even more closely matched with a 12 per cent discrepancy. The disagreement between the two methods can be explained by the relative weighting of warm high-column density gas with bright 13CO (3–2) emission and colder low-column density gas with that is faint in 13CO (3–2) emission in the determination of a single value.
@@
For both regions, the relationship between H2 column density and integrated intensity of 12CO and 13CO (3–2) emission is, of course, more complex than a simple multiplicative factor as can clearly be seen in Fig. 7 panels a) and d). Barnes et al. (2015) remarked on similar behaviour for 12CO (1–0) emission and suggested adopting a power law relationship for the column density of CO where , however the relationships that we report are clearly also not linear in log-space indicating that even a single power law would not accurately reproduce the behaviour. A power-law with an index of would help reproduce the super-linear parts of the distributions in Figures 7 a) and d), but these would then under-predict (H2) at low CO intensity. Our method ii) values that incorporate a background column density could help alleviate this issue, but we can not be certain whether the background column density is a result of the greater column of Galactic disc that the Inner Galaxy sight-line contains, or if it is reflective of a difference in excitation conditions of CO; sub-thermal emission of CO would also cause an flattening of the relationship at low CO intensity. We therefore recommend that the single representative values should be used in the knowledge that more accurate column densities may be determined through LTE modelling of the combination of 12CO and 13CO emission in combination, where available. We will perform the a comparison of LTE-derived excitation conditions between the Inner Galaxy and Outer Galaxy data in a future paper.
@@
In Section 4.2, we also compared the properties of sources extracted from PAMS to the those from a different Inner Galaxy sample from CHIMPS. Here, we created a distance-limited sample of CHIMPS clouds with distances between 2–4 kpc – approximately matching the PAMS range of distances – corresponding to Galactocentric radii of 5–7 kpc. In Fig. 8) we found no substantial differences between two types of radius measurement between the two samples, when considering a source extractions designed to be as consistent as possible, indicating that the emission profiles of sources in the two samples are similar. The size–linewidth and mass–radius relationships of the two samples of extracted sources sources in Fig. 9 were again found to be generally consistent.
One notable difference is that the masses of PAMS sources tend to be greater for a given size-scale, with difference between the distributions of of 0.7 dex. This effect also manifested in Fig. 10 where the mean density of PAMS sources at a kpc is similarly elevated by dex compared to sources with kpc. Some of this can be accounted for by the fact that the Herschel column densities we used – determined by Marsh et al. (2017) to derive our factors used for the mass determinations (Marsh et al., 2017) adopted a single gas-to-dust mass ratio which we know has a Galactocentric dependence (Giannetti et al., 2017). The Giannetti et al. (2017) relation gives a good match with the gas-to-dust mass ratio of 100 used by Marsh et al. (2017) for sources at 6 kpc, but the value is a factor of 2 higher at 9.5 kpc, accounting for 0.3 dex of the above discrepancies. Selection biases resulting from the PAMS survey being targeted towards well-known star-forming regions (in contrast to the blind survey mode of CHIMPS) may account for part of the remaining discrepancy. Finally we also used the same value of for the mass determinations was the same in the Outer and Inner Galaxy following our method i) values in Section 4.1, but both methods iii) and iv) suggest a larger value for the Inner Galaxy which could also account for a further 0.2–0.3 dex of discrepancy. Overall, we do not find any significant differences between the sources in the Inner and Outer Galaxy over kpc.
@@ 6 Conclusions
We have presented data from the Perseus Arm Molecular Survey (PAMS) in 13CO and C18O (3–2), a survey of 6 deg2 covering several molecular cloud complexes in the Outer Galaxy, at Galactocentric radii of 9.5 kpc. We incorporated archival 12CO (3–2) data, which cover some of the PAMS regions fully, and others only partially. In Section 4.1, we calculated factors for 12CO and 13CO (3–2), which convert the integrated intensity of the CO emission into molecular hydrogen column density, and examined the distributions of pixel values. In Sections 4.2.1, 4.2.2, and 4.3, we compared basic properties of sources extracted from PAMS with equivalent values from the CHIMPS survey in order to probe different Galactic environments.
@@
The systematic variation in values derived the Inner and Outer Galaxies were generally small compared to the variation arising from the different methods, as well as the spread in individual pixel values, but that the difference was stronger in 13CO and 12CO.
@@
We recommend the usage of a value of = cm-2 (km s-1)-1 and cm-2 (km s-1)-1 for resolved and unresolved applications, respectively, to convert integrated intensity of 12CO (3–2) emission to molecular hydrogen column density.
@@
For 13CO (3–2), we recommend the usage of a value of = cm-2 (km s-1)-1 and cm-2 (km s-1)-1 for resolved and unresolved applications, respectively, to convert to molecular hydrogen column density.
@@
5.
Although the distributions masses of PAMS Outer Galaxy sources are shifted to greater values than their CHIMPS Inner Galaxy counterparts, most of the 0.6 dex difference can be explained by the Galactic gradient in gas-to-dust mass ratios and selection biases.
Finally, we have demonstrated that the PAMS data are a valuable addition to the existing repertoire of publicly-available 12CO, 13CO, and C18O (3–2) survey data in the Outer Galaxy. In combination with surveys such as CHIMPS, CHIMPS2, and COHRS, these data extend the baseline in Galactocentric radius in what can be studied in the 3–2 rotational transition across different Galactic environments, in addition to the growing number of surveys covering 2–1 and 1–0.
Acknowledgments
AJR acknowledges postdoctoral support from the University of Leeds. The JCMT is operated by the East Asian Observatory on behalf of The National Astronomical Observatory of Japan; Academia Sinica Institute of Astronomy and Astrophysics; the Korea Astronomy and Space Science Institute; the National Astronomical Research Institute of Thailand; Center for Astronomical Mega-Science (as well as the National Key R&D Program of China with No. 2017YFA0402700). Additional funding support is provided by the Science and Technology Facilities Council of the United Kingdom and participating universities and organizations in the United Kingdom and Canada. The JCMT has historically been operated by the Joint Astronomy Centre on behalf of the Science and Technology Facilities Council of the United Kingdom, the National Research Council of Canada and the Netherlands Organisation for Scientific Research. This research used the facilities of the Canadian Astronomy Data Centre operated by the National Research Council of Canada with the support of the Canadian Space Agency. This research has made use of NASA’s Astrophysics Data System Bibliographic Services’
@@ Acknowledgments
We make the PAMS data publicly available at https://www.canfar.net/storage/vault/list/PAMS (will be replaced with DOI when finalised). This repository includes the 13CO and C18O (3–2) mosaics of each region, along with the 12CO (3–2) cubes produced from the raw time-ordered data from the JCMT archive where available. The repository also includes the FellWalker catalogues and the corresponding assignment masks.
@@ Appendix A ORAC-DR Recipe Parameters
In this Appendix, we list of several sections of the recipe parameters used for the data reduction recipe reduce_science_narrowline in orac-dr. The first block relates to the pixel size and the binning method:
@@
Which specifies the use of 6.0 arcsecond-wide pixels, and the use of a Gaussian smoothing kernel with FWHM of 8-arcseconds to assist with the binning of pixel values onto the new pixel grid. The spreading function is curtailed at 6-arcseconds, as specified by SPREAD_FWHM_OR_ZERO. Finally, the REBIN parameter specifies that the cube will be regridded to a 0.3 km s-1-wide velocity channels.
@@
The above parameters specify that a 1st order polynomial baseline should be used for the fitting, and that the region of 80 to km s-1 in the spectrum should be excluded when performing baseline linearity tests for each receptor.
@@
This final set of parameters was optionally used in instances where off-position absorption was suspected to be present. This often reveals itself as a velocity range showing absorption features in the cube-average spectrum. By enabling SUBTRACT_REF_SPECTRUM and REF_SPECTRUM_COMBNINE_REFPOS, orac-dr interpolates the reference spectrum across the velocity range or ranges identified in the REF_SPECTRUM_REGIONS setting, which consists of a comma-separated list of regions with suspected off-position emission. In this particular instance, the ranges of 15.0 to and 14.0 to 17.0 km s-1 were interpolated over in the off- (reference-) position spectrum.
Appendix B FellWalker configuration
In Section 3.3 we described our usage of FellWalker to extract sources from the 13CO (3–2) PAMS data. Our parameter setup for the fwhires extraction was:
@@ Appendix B FellWalker configuration
The parameter selection for the fwlores extraction was identical, with the exception of:
@@
This change effectively suppresses the ability of FellWalker to identify substructures within isolated islands of emission, and will this recover the largest possible complexes of contiguous pixels of emission.
The parameters used for the fwchimps extraction were the same as for fwhires with the following alterations:
@@
This setup was chosen to match the extraction of Rigby et al. (2016) as closely as possible. The key difference is that was RMS is set to 1.7 to reflect that higher noise levels in the CHIMPS compared to PAMS data after smoothing to 27.4 arcsec resolution. The parameters FlatSlope, MinDip, MinHeight, and Noise are scaled up by the same amount. The differences between the fwhires setup compared to fwchimps were generally selected to combat spurious sources, and especially the archipelago source types discussed briefly in Section 4.2; setting both CleanIter and MaxJump to zero, and VeloRes to 1 helps suppress these source types. It was also necessary to adopt value of 1 for AllowEdge due to the extra sensitivity and targeted mapping of the PAMS observations, which meant that emission was often detected up to the edges of the image, and a zero value would have excluded significant sources.

### v3

@@ 5.1 Galactocentric gradients in
In Section 4.1, we calculated factors for 12CO and 13CO (3–2) between the Inner and Outer Galaxy by comparing Herschel-derived H2 column densities with data from the COHRS (Dempsey et al., 2013; Park et al., 2023), CHIMPS (Rigby et al., 2016), and PAMS data, summarising the results in Tables 4 and 4. The Inner Galaxy sight-line used was at °, whose emission is dominated by the W43 star-forming region at a distance of 5.2 kpc (Urquhart et al., 2018), corresponding to a Galactocentric radius of kpc. We assume that our Inner Galaxy -factors are weighted somewhere slightly closer along the line of sight than W43, so we will take kpc as the representative value for the Inner Galaxy results777The intensity-weighted Galactocentric radius from the Rigby et al. (2019) catalogue of clumps in the longitude range of the tile gives a compatible result., while the PAMS regions are located close to kpc.
We found that the distributions of and values calculated on a pixel-by-pixel basis were significantly different between the Inner and Outer Galaxy regions (and this is confirmed by both Kolmogorov-Smirnov and Anderson-Darling statistical tests). For 12CO, the difference is relatively small. Our suggested global value of = cm-2 (K km s-1)-1 is weighted towards the Inner Galaxy value, and is identical to the value adopted by Colombo et al. (2019) who also used COHRS data. If one considers a common formulation of
where is the 12CO (3–2)/(1–0) line ratio, these results imply an average value of , which is consistent with observations of nearby galaxies (e.g Lamperti et al., 2020), high-redshift galaxies (e.g. Aravena et al., 2014) and simulations (e.g. Peñaloza et al., 2018). The Inner and Outer Galaxy distributions of values in Fig. 7 overlap substantially, and representative values for the Outer Galaxy tend to be a factor of 1.3–2.0 lower than their Inner Galaxy counterparts (see Table 4). Taking equation (12) literally, these results would imply that increases with Galactocentric distance, from 0.5 to 0.7 between 5.0 and 9.5 kpc based on equation (refeq:R31). This is contrary to the shallow gradients in (which we assume would hold the same trend) found by Sakamoto et al. (1997) in the Milky Way and Den Brok et al. (2022) in M51, but this implication would only hold if the values of and do not vary at the same time. Since we are not directly probing the intensity ratio here, other factors may affect this result.
In equation (4) in Section 4.1 we used a single value of for the gas-to-dust mass ratio, which was shown by Giannetti et al. (2017) to have a Galactocentric dependence (with a large scatter). The Galactic gradient presented in that study would imply a variation of to 185 between 5.0 to 9.5 kpc, and such a variation could reverse the apparent increasing trend in to a decreasing one, and simultaneously reduce the magnitude of the trend. Rescaling the Inner and Outer Galaxy values based on this gradient would result in and cm-2 (K km s-1)-1 for the Inner and Outer Galaxy, respectively, implying decreasing from 0.65 to 0.45 over the same range. For 13CO (3–2), the trend also reverses, resulting in and cm-2 (K km s-1)-1 for the Inner and Outer Galaxy, respectively. Considering this variation, a value of cm-2 (K km s-1)-1 is appropriate for general use in the Galactic disc for sources lying within 4–10 kpc of the Galactic Centre, and incorporates the gradient. A factor of 1.5 uncertainty encapsulates much of the variation we have seen.
Our Inner and Outer Galaxy values of and do not vary significantly, given the uncertainties in the gas-to-dust mass ratio, and our results are not inconsistent with a radially decreasing . More-robust direct measurements of could be obtained in future at 50-arcseconds resolution using the MWISP data, or at 20-arcsec resolution using a combination of CHIMPS, PAMS, COHRS, and FUGIN data.
@@
In all cases, the relationship between H2 column density and integrated intensity of 12CO and 13CO (3–2) emission is, of course, more complex than a simple multiplicative factor as can clearly be seen in Fig. 8 Panels a) and d). Barnes et al. (2015) remarked on similar behaviour for 12CO (1–0) emission and suggested adopting a power-law relationship for the column density of CO where , however, the relationships that we report are clearly also not linear in log-space, indicating that even a single power law would not accurately reproduce the behaviour. A power law with an index of would help reproduce the super-linear parts of the distributions in Figs. 8 a) and d), but these would then under-predict (H2) at low CO intensity. Our Method ii) values that incorporate a background column density could help alleviate this issue, but we cannot be certain whether the background column density is a result of the greater column of Galactic disc that the Inner Galaxy sight-line contains, or if it is reflective of a difference in excitation conditions of CO; sub-thermal emission of CO would also cause an flattening of the relationship at low CO intensity. We therefore recommend that the single representative values should be used in the knowledge that more-accurate column densities may be determined through LTE modelling of the combination of 12CO, 13CO, and – where available – C18O emission in combination.
5.2 Galactocentric gradients in molecular-cloud properties
@@
In Section 4.2, we compared the properties of sources extracted from PAMS with those from a different Inner Galaxy sample from CHIMPS. Here, we created a distance-limited sample of CHIMPS clouds with distances between 2–4 kpc – approximately matching the PAMS range of distances – corresponding to Galactocentric radii of 5–7 kpc. In Fig. 9 we found no substantial differences between two types of radius measurement between the two samples, when considering the fwchimps sources extraction that was designed to be replicate the CHIMPS setup as closely as possible, indicating that the emission profiles of sources in the two samples are similar. The size–linewidth and mass–radius relationships of the two samples of extracted sources sources in Fig. 10 were again found to be consistent.
One notable difference is that the masses of PAMS sources tend to be greater for a given size-scale (Fig. 10), with difference between the distributions of of 0.15 dex (i.e. around 40 per cent). A surprising result follows, which is that a greater proportion of the PAMS sample satisfy the Kauffmann & Pillai (2010) criterion for HMSF than for CHIMPS, with 96 and 63 per cent of the fitted samples doing so, respectively. At face value, this would suggest that the molecular clouds at kpc have more massive star-forming potential than those at kpc, which is contrary to our expectations, e.g. the decreasing fraction of star-forming Hi-GAL clumps reported by Ragan et al. (2016). Since the fwchimps source extraction accounts for the differences in data quality between PAMS and CHIMPS, we can rule out biases arising from differences in sensitivity. This discrepancy is made worse if we account for the Galactic gradient in the gas-to-dust mass ratio , discussed in the previous section, which would suggest that our Inner and Outer Galaxy factors were over- and under-estimated, respectively. On the other hand, Giannetti et al. (2017) point to large uncertainties on for individual star-forming regions, as well as intrinsic scatter, and so this may partly be a result of the small number of independent star-forming complexes that PAMS covers. Additionally, while the Outer Galaxy values derived in Section 4.1 for the various methods varied little, the Inner Galaxy values were wider-ranging, and the Method iii) value is 50 per cent larger than the Method i) value we adopted, and this may account for another part of the discrepancy. Furthermore, selection biases resulting from the PAMS survey being targeted towards well-known star-forming regions (in contrast to the blind survey mode of CHIMPS) may account for a further part of the discrepancy. While we can find no single explanation that is entirely satisfactory, we suggest a combination of the above factors may explain this result. Overall, we do not find any significant differences between the sources in the Inner and Outer Galaxy over kpc.
@@ 6 Conclusions
We have presented the Perseus Arm Molecular Survey (PAMS), a survey of 12CO, 13CO, and C18O (3–2), covering 8 deg2 over several molecular-cloud complexes in the Outer Galaxy, at Galactocentric radii of 9.5 kpc. In Section 4.1, we calculated factors for 12CO and 13CO (3–2), which convert the integrated intensity of the CO emission into molecular-hydrogen column density, and examined the distributions of pixel values. In Section 4.2, we compared basic properties of sources extracted from PAMS with equivalent values from the CHIMPS survey in order to probe any differences between these Inner and Outer Galaxy environments, covering a range of Galactocentric radii between 4–10 kpc.
@@
The systematic variation in values derived the for Inner and Outer Galaxy were generally small compared with the variation arising from the different methods, as well as the spread in individual pixel values, but that difference was stronger in 13CO (3–2) than 12CO (3–2).
@@
We recommend the usage of a value of = cm-2 (K km s-1)-1 to convert 12CO (3–2) integrated intensity to molecular-hydrogen column density, with a factor of 1.5 uncertainty, in agreement with previous studies (e.g. Colombo et al., 2019).
@@
For 13CO (3–2), we recommend the usage of a value of cm-2 (K km s-1)-1 to convert 13CO (3–2) integrated intensity to molecular-hydrogen column density, with a factor of 1.5 uncertainty.
@@
Although the values we recovered are significantly different for the Inner and Outer Galaxy (traced by CHIMPS and PAMS), with cm-2 (K km s-1)-1 for the Inner Galaxy and cm-2 (K km s-1)-1 for the Outer Galaxy, accounting for the Galactic gradient in gas-to-dust mass ratio can resolve much of the difference. Calculating H2 column densities from LTE analysis of co-spatial 12CO and 13CO emission would be preferable to relying on factors.
5.
@@
6.
Although the distributions of masses of PAMS Outer Galaxy sources are shifted to greater values than their CHIMPS Inner Galaxy counterparts for a given size scale, most of the 0.2 dex difference can be explained by a combination of variations in gas-to-dust mass ratios and selection biases.
Finally, we have demonstrated that the PAMS data are a valuable addition to the existing repertoire of publicly available 12CO, 13CO, and C18O (3–2) survey data in the Outer Galaxy. In combination with surveys such as CHIMPS, CHIMPS2, and COHRS, these data extend the baseline in Galactocentric radius in what can be studied in the 3–2 rotational transition across different Galactic environments, in addition to the growing number of surveys covering 2–1 and 1–0.
Acknowledgements
We would like to thank the anonymous referee whose review helped to improve the quality and accuracy of this manuscript. We thank Andrea Giannetti for a thoughtful discussion about gas-to-dust mass ratios. AJR acknowledges postdoctoral support from the University of Leeds. The JCMT is operated by the East Asian Observatory on behalf of The National Astronomical Observatory of Japan; Academia Sinica Institute of Astronomy and Astrophysics; the Korea Astronomy and Space Science Institute; the National Astronomical Research Institute of Thailand; Center for Astronomical Mega-Science (as well as the National Key R&D Program of China with No. 2017YFA0402700). Additional funding support is provided by the Science and Technology Facilities Council of the United Kingdom and participating universities and organizations in the United Kingdom and Canada. The JCMT has historically been operated by the Joint Astronomy Centre on behalf of the Science and Technology Facilities Council of the United Kingdom, the National Research Council of Canada and the Netherlands Organisation for Scientific Research. The authors wish to recognize and acknowledge the very significant cultural role and reverence that the summit of Maunakea has always had within the indigenous Hawaiian community. We are most fortunate to have the opportunity to conduct observations from this mountain. This research used the facilities of the Canadian Astronomy Data Centre operated by the National Research Council of Canada with the support of the Canadian Space Agency. This research has made use of NASA’s Astrophysics Data System Bibliographic Services’.
Software
Python packages:
astropy (The Astropy Collaboration, 2022),
ipython (Perez & Granger, 2007),
matplotlib (Hunter, 2007),
multicolorfits (Cigan, 2019),
numpy (Harris et al., 2020),
scipy (Virtanen et al., 2020),
scikit-learn (Pedregosa et al., 2011).
@@ Acknowledgements
We make the PAMS data publicly available at https://dx.doi.org/10.11570/25.0001. This repository includes the 12CO, 13CO, and C18O (3–2) mosaics of each region, and the FellWalker catalogues with their corresponding assignment masks.
@@ Appendix A ORAC-DR Recipe Parameters
In this Appendix, we list of several sections of the recipe parameters used for the data-reduction recipe reduce_science_narrowline in orac-dr. The first block relates to the pixel size and the binning method:
@@
which specifies the use of 6.0-arcsecond-wide pixels, and the use of a Gaussian smoothing kernel with FWHM of 8 arcsec to assist with the binning of pixel values onto the new pixel grid. The spreading function is curtailed at 6 arcsec, as specified by SPREAD_FWHM_OR_ZERO. Finally, the REBIN parameter specifies that the cube will be regridded to a 0.3 km s-1-wide velocity channels.
@@
The above parameters specify that a first-order polynomial baseline should be used for the fitting, and that the region of 80 to km s-1 in the spectrum should be excluded when performing baseline linearity tests for each receptor.
@@
This final set of parameters was optionally used in instances where off-position absorption was suspected to be present. This often reveals itself as a velocity range showing absorption features in the cube-average spectrum. By enabling SUBTRACT_REF_SPECTRUM and REF_SPECTRUM_COMBINE_REFPOS, orac-dr interpolates the reference spectrum across the velocity range or ranges identified in the REF_SPECTRUM_REGIONS setting, which consists of a comma-separated list of regions with suspected off-position emission. In this particular instance, the ranges of 15.0 to and 14.0 to 17.0 km s-1 were interpolated over in the off- (reference-) position spectrum.
Appendix B Preview images of the PAMS regions
In this appendix we display integrated intensity (moment 0) images of the PAMS regions in 12CO, 13CO, and C18O(3–2), along with the corresponding position-velocity diagrams. We do not include NGC 7538 which has already been shown in Fig. 5. In all cases, the cubes have been masked using FellWalker, adopting the fwlores masks discussed in Section 3.3.
Appendix C FellWalker configuration
In Section 3.3 we described our usage of FellWalker to extract sources from the 13CO (3–2) PAMS data. Our parameter setup for the fwhires extraction was as follows.
@@ Appendix C FellWalker configuration
The parameter selection for the fwlores extraction was identical, with the following exception.
@@
This change effectively suppresses the ability of FellWalker to identify substructures within isolated islands of emission, and thus will recover the largest-possible complexes of contiguous pixels of emission.
The parameters used for the fwchimps extraction were the same as for fwhires with the following alterations.
@@
This setup was chosen to match the extraction of Rigby et al. (2016) as closely as possible. The key difference is that RMS is set to 1.7 to reflect higher noise levels in the CHIMPS compared with PAMS data after smoothing to 27.4-arcsec resolution. The parameters FlatSlope, MinDip, MinHeight, and Noise are scaled up by the same amount. The differences between the fwhires setup compared with fwchimps were generally selected to combat spurious sources, and especially the archipelago source types discussed briefly in Section 4.2; setting both CleanIter and MaxJump to zero, and VeloRes to 1 helps suppress these source types. It was also necessary to adopt value of 1 for AllowEdge due to the extra sensitivity and targeted mapping of the PAMS observations, which meant that emission was often detected up to the edges of the image, and a zero value would have excluded significant sources.
Appendix D FellWalker catalogue descriptions
The columns featuring in the fwhires, fwlores, and fwchimps catalogues are listed Table 6. The only difference between the three is that the fwhires catalogue features an additional Parent_ID column, which allows the fwhires sources to be connected to their parent larger-scale structure in the fwlores extraction.
| Column number | Name | Data type | Units | Description |
| 1 | Region |
string | None | PAMS region |
| 2 | ID |
integer | None | Voxel ID in mask |
| 3 | RA_cen |
float | deg | Centroid RA coordinate |
| 4 | Dec_cen |
float | deg | Centroid Dec coordinate |
| 5 | Vlsr_cen |
float | km s-1 | Centroid coordinate |
| 6 | RA_peak |
float | deg | Peak RA coordinate |
| 7 | Dec_peak |
float | deg | Peak Dec coordinate |
| 8 | Vlsr_peak |
float | km s-1 | Peak coordinate |
| 9 | Sigma_RA |
float | arcsec | Intensity-weighted dispersion in RA |
| 10 | Sigma_Dec |
float | arcsec | Intensity-weighted dispersion in Dec |
| 11 | Sigma_Vlsr |
float | km s-1 | Intensity-weighted dispersion in |
| 12 | Sigma_Vlsr_corrected |
float | km s-1 | Estimated linewidth corrected for SNR |
| 13 | Sigma_Vlsr_err |
float | km s-1 | Uncertainty on linewidth |
| 14 | Sum |
float | K | Sum of 13CO (3–2) voxel values |
| 15 | Peak |
float | K | Peak 13CO (3–2) voxel value |
| 16 | major_sigma |
float | deg | Semi-major axis of ellipse |
| 17 | minor_sigma |
float | deg | Semi-minor axis of ellipse |
| 18 | position_angle |
float | deg | Position angle of ellipse (E of N) |
| 19 | n_pixels |
integer | None | Number of pixels in projected area |
| 20 | n_voxels |
integer | None | Number of voxels in source |
| 21 | polygon |
string | None | (RA, Dec) vertices of polygon describing source |
| 22 | RA_cen_snr |
float | deg | Centroid RA coordinate in SNR cube |
| 23 | Dec_cen_snr |
float | deg | Centroid Dec coordinate in SNR cube |
| 24 | Vlsr_cen_snr |
float | km s-1 | Centroid coordinate in SNR cube |
| 25 | RA_peak_snr |
float | deg | Peak RA coordinate in SNR cube |
| 26 | Dec_peak_snr |
float | deg | Peak Dec coordinate in SNR cube |
| 27 | Vlsr_peak_snr |
float | km s-1 | Peak coordinate in SNR cube |
| 28 | Sigma_RA_snr |
float | arcsec | Intensity-weighted dispersion in RA in SNR cube |
| 29 | Sigma_Dec_snr |
float | arcsec | Intensity-weighted dispersion in Dec in SNR cube |
| 30 | Sigma_Vlsr_snr |
float | km s-1 | Intensity-weighted dispersion of in SNR cube |
| 31 | Sum_snr |
float | None | Sum of voxel values in SNR cube |
| 32 | Peak_snr |
float | None | Peak voxel value in SNR cube |
| 33 | dkpc |
float | kpc | Heliocentric distance to region |
| 34 | derr |
float | kpc | Error on heliocentric distance to region |
| 35 | Rgc |
float | kpc | Galactocentric radius of region |
| 36 | R_eq_arcsec_nodec |
float | arcsec | Radius of circle with equivalent area |
| 37 | R_eq_arcsec |
float | arcsec | Beam-deconvolved radius of circle with equivalent area |
| 38 | R_eq_pc |
float | pc | Beam-deconvolved radius of circle with equivalent area |
| 39 | R_eq_pc_err |
float | pc | Error on R_eq_pc
| 40 | R_sig_arcsec_nodec |
float | arcsec | Intensity-weighted radius |
| 41 | R_sig_arcsec |
float | arcsec | Beam-deconvolved intensity-weighted radius |
| 42 | R_sig_pc |
float | pc | Beam-deconvolved intensity-weighted radius |
| 43 | R_sig_pc_err |
float | pc | Error on R_sig_pc
| 44 | Mass_Xco |
float | M⊙ | Mass estimated using 13CO (3–2) X-factor |
| 45 | Mass_Xco_lo |
float | M⊙ | 16th percentile of Mass_Xco distribution |
| 46 | Mass_Xco_up |
float | M⊙ | 84th percentile of Mass_Xco distribution |
| 47* | Parent_ID |
integer | None | ID of parent source in fwlores catalogue |
* (fwhires catalogue only) |